Hi, I'm very stuck on this question! The answers 20.47 and 26.02 were wrong. Thanks!

A 0.0500-kg ice cube at −30.0°C is placed in 0.500 kg of 35.0°C water in a very well-insulated container. What is the final temperature in degrees Celsius?
Answer #1

When the ice and the water are mixed, first temperaturevof ice rises to its melting point of 0°C , then melting of the ice starts at constant temperature. Once whole ice is molten, temperature of this water form rises till equilibrium temperature. At the equilibrium:

Heat gain by ice = Heat loss by the water.

5080.5830 + SoX 80 + 50XIXT= soox 1x (35-T) 15+ 80 + T = 10(35-T) -> IT 350-95 T = 255 = 23.18 °C